# Generative design of functional organic molecules for terahertz radiation detection

### v1

@@ III.2 Evaluation and improvement of property predictors
As shown above, generated molecules significantly differ in chemical composition from the training molecules. This raises the question of whether the KRR predictor of the THz upconversion efficiency metric, P, provides transferable prediction accuracy for the novel, generated molecules – a crucial prerequisite for targeted property-driven molecular design. To assess this, DFT structure optimisations and vibrational spectrum calculations were performed on randomly selected molecules from the Thiol database that was used to train the G-SchNet model and from the dataset generated in Iteration 6. Table 1 shows the performance of the KRR predictor on these molecules. The mean absolute error (MAE) on the Thiol database is similar to the MAE on the test set of the THz database, while the MAE increases significantly for molecules generated in Iteration 6. In particular, the KRR model severely underestimates the P values of high-P molecules, as shown in the ESI (Figure S6), which suggests that the true P values of molecules generated in the biasing workflow reach much higher values than what is predicted in Fig. 1a.
@@
As the KRR predictor uses SMILES strings as input and is based on two-dimensional Morgan fingerprints, it does not benefit from the information contained in the 3D structures generated by G-SchNet. As the THz upconversion efficiency sensitively depends on the molecular conformation and vibrational frequencies, this limits the expressiveness and prediction accuracy of the model. We therefore trained two equivariant GNN models with three-dimensional atom-wise embeddings on the same THz dataset, namely the MACE and PaiNN models. Table 1 compares the MAE of the different ML models for the reference DFT-calculated P values, determined for the DFT-optimized structures of test molecules from the THz dataset. Both MACE and PaiNN provide improved predictions compared to the EN and KRR models of Koczor-Benda et al. [10], with PaiNN providing the best prediction. PaiNN also learns faster than MACE from less data, as shown by the learning curves in the ESI (Fig. S3); for this reason, the PaiNN predictor was used for all subsequent analyses. When testing the PaiNN model on the molecules generated in Iteration 6, the MAE is larger with 0.73 (Table 1). PaiNN also underestimates the P values of high-P value molecules, as shown in the ESI (Fig. S6), though this is slightly less pronounced than with KRR. Therefore, all tested models show reduced prediction accuracy when applied to the iteratively biased datasets, suggesting that the models are forced to predict outside of the chemical space spanned by the training data. This severely limits their ability to act as a transferable property predictor that drives molecule generation. The deterioration of the model accuracy for the THz upconversion efficiency is more significant than what was observed by Westermayr et al. [16] for electronic property prediction. We hypothesize that this is due to the integrated nature of the THz upconversion metric P and its sensitive dependence on collective low-frequency molecular vibrations and the molecular polarizability.
@@ III.4 Identification of candidate molecules
We selected generated molecules with (based on predictions by the retrained PaiNN predictor) and employed AiZynthFinder to perform retrosynthetic planning. From the 1011 molecules satisfying the selection criterion, only 34 were predicted to have retrosynthetic routes from purchasable precursors [52] based on a stock from compounds available within the ZINC [53] database; retrosynthetic paths for these molecules can be found in Fig. S11–S17. Notably, all 34 molecules belong to clusters from which subclusters C2 and C5 were drawn (ESI, Table SVI).
@@
To confirm the suitability of these molecules for THz radiation detection, their absorption, Raman scattering and frequency upconversion spectra were calculated, and their P values were determined using DFT. The top candidate, visualized in Fig. 4, has a DFT-calculated P value of 7.88. Considering that the P value is a logarithmic quantity (equation 1), this is significantly higher than any of the molecules previously identified within commercial databases in Koczor-Benda et al. [10], where the top 5 candidates had P values between 5.30 and 6.18. Figure 4 also shows the relevant properties and vibrational spectra of the top candidate, while vibrational spectra and properties of other candidate molecules with DFT-calculated P values above 5.20 are shown in the ESI (Fig. S18–S21). The top molecule has two vibrational modes that are highly active in frequency upconversion, which are located at 515 and 559 . Both modes involve an out-of-plane (umbrella) motion of one of the amino groups that is coupled to out-of-plane motions of hydrogen atoms of the neighboring ring. This out-of-plane motion of the amino group is also responsible for the highest intensity peaks of other top candidates, as shown in the ESI (Fig. S18–S21). This provides evidence that the aromatic amine functional group not only correlates with high P values, but is also directly involved in the upconversion process. The highly active mode appears in the 515–832 spectral range for the top candidates, showing that the chemical environment and the coupling of the out-of-plane motion of the amino group with other vibrations of the molecule have a significant effect on the position of the peak. This can be advantageous for the tuning of narrowband THz radiation detectors operating at different frequencies. We also note that within the top candidates, molecules with the same SMILES string were generated multiple times with different 3D structures in the different biasing iterations. As the SCScore and KRR-predicted P values depend only on 2D information, they remain the same for different conformers. However, the PaiNN-predicted P values for raw generated structures and DFT-calculated P values for structures that have undergone geometry optimization can differ, as shown in section S10 and Fig. S23 of the ESI. This further highlights the benefits of working with property predictors that are based on 3D descriptors.
@@
Of the 34 molecules listed in Table SVI, only one compound (generated three times as different conformers, all sharing the same SMILES string) was identified in the PubChem [56, 57] database, Nc1cc(S)c(cc1N)N, which corresponds to 2,4,5-triaminobenzenethiol (Compound Identifier 67981805 [62]). The remaining 31 molecules were not found in PubChem, likely representing novel candidate structures THz upconversion applications.
@@ IV Conclusions and Outlook
V Data availability
Data for this article, including molecular databases in ASE database format, DFT-optimized best candidate molecules, and ASE databases for xTB calculations are available online: https://doi.org/10.6084/m9.figshare.28539995.v1 [63]. Code for the extraction of bonding features from molecular databases and obtaining the principal components of the structural/bonding descriptors has been released in our GSchNetTools package, available at https://github.com/maurergroup/GSchNetTools.
VI Acknowledgements

### v2

@@ III.2 Evaluation and improvement of property predictors
As shown above, generated molecules significantly differ in chemical composition from the training molecules. This raises the question of whether the KRR predictor of the THz upconversion efficiency metric, P, provides transferable prediction accuracy for the novel, generated molecules – a crucial prerequisite for targeted property-driven molecular design. To assess this, DFT structure optimizations and vibrational spectrum calculations were performed on randomly selected molecules from the Thiol database that was used to train the G-SchNet model and from the dataset generated in Iteration 6. Table 1 shows the performance of the KRR predictor on these molecules. The mean absolute error (MAE) on the Thiol database is similar to the MAE on the test set of the THz database, while the MAE increases significantly for molecules generated in Iteration 6. In particular, the KRR model severely underestimates the P values of high-P molecules, as shown in the ESI (Figure S6), which suggests that the true P values of molecules generated in the biasing workflow reach much higher values than what is predicted in Fig. 1a.
@@
As the KRR predictor uses SMILES strings as input and is based on 2D Morgan fingerprints, it does not benefit from the information contained in the 3D structures generated by G-SchNet. As the THz upconversion efficiency sensitively depends on the molecular conformation and vibrational frequencies, this limits the expressiveness and prediction accuracy of the model. We therefore trained two equivariant GNN models with 3D atom-wise embeddings on the same THz dataset, namely the MACE and PaiNN models. Table 1 compares the MAE of the different ML models for the reference DFT-calculated P values, determined for the DFT-optimized structures of test molecules from the THz dataset. Both MACE and PaiNN provide improved predictions compared to the EN and KRR models of Koczor-Benda et al. [11], with PaiNN providing the best prediction. PaiNN also learns faster than MACE from less data, as shown by the learning curves in the ESI (Fig. S3); for this reason, the PaiNN predictor was used for all subsequent analyses. When testing the PaiNN model on the molecules generated in Iteration 6, the MAE is larger with 0.73 (Table 1). PaiNN also underestimates the P values of high-P value molecules, as shown in the ESI (Fig. S6), though this is slightly less pronounced than with KRR. Therefore, all tested models show reduced prediction accuracy when applied to the iteratively biased datasets, suggesting that the models are forced to predict outside of the chemical space spanned by the training data. This severely limits their ability to act as a transferable property predictor that drives molecule generation. The deterioration of the model accuracy for the THz upconversion efficiency is more significant than what was observed by Westermayr et al. [17] for electronic property prediction. We hypothesize that this is due to the integrated nature of the THz upconversion metric P and its sensitive dependence on collective low-frequency molecular vibrations and the molecular polarizability.
@@ III.4 Identification of candidate molecules
We selected generated molecules with (based on predictions by the retrained PaiNN predictor) and employed AiZynthFinder to perform retrosynthetic planning. From the 1011 molecules satisfying this selection criterion, only 34 were predicted to have retrosynthetic routes from purchasable precursors [56] based on a stock from compounds available within the ZINC [57] database; retrosynthetic paths for these molecules can be found in Fig. S11–S17. Notably, all 34 molecules belong to clusters from which subclusters C2 and C5 were drawn (ESI, Table SVI).
To confirm the suitability of these molecules for THz radiation detection, their absorption, Raman scattering and frequency upconversion spectra were calculated, and their P values were determined using DFT. Figure 4 shows the relevant properties and vibrational spectra of the top candidate, while vibrational spectra and properties of other candidate molecules with DFT-calculated P values above 5.20 are shown in the ESI (Fig. S18–S21). The top candidate, 2-amino-5-(4-aminophenylamino)pyridine-4-thiol has a DFT-calculated P value of 7.88. Considering that the P value is a logarithmic quantity (Equation (1)), this is significantly higher than any of the molecules previously identified within commercial databases in Koczor-Benda et al. [11], where the top 5 candidates had P values between 5.30 and 6.18. For the fifth top molecule from Koczor-Benda et al. [11], 5-amino-2-mercaptobenzimidazole, Redolat et al. [5] developed a functionalization technique to prepare self-assembled molecular monolayers in gold-based plasmonic nanocavities and successfully integrated these nanocavities on a silicon-based photonic chip. While frequency upconversion measurements are not yet available for this compound, our DFT simulations suggest about 14 times higher upconversion capability in the THz/mid-infrared range for the most active mode (559 cm-1) of our top candidate compared to the most active mode (458 cm-1) of 5-amino-2-mercaptobenzimidazole (see Fig. S22 in ESI).
@@
The top molecule has two vibrational modes that are highly active in frequency upconversion, which are located at 515 and 559 . Both modes involve an out-of-plane (umbrella) motion of one of the amino groups that is coupled to out-of-plane motions of hydrogen atoms of the neighboring ring. This out-of-plane motion of the amino group is also responsible for the highest intensity peaks of other top candidates, as shown in the ESI (Fig. S18–S21). This provides evidence that the aromatic amine functional group not only correlates with high P values, but is also directly involved in the upconversion process. The highly active mode appears in the 515–832 spectral range for the top candidates, showing that the chemical environment and the coupling of the out-of-plane motion of the amino group with other vibrations of the molecule have a significant effect on the position of the peak. This can be advantageous for the tuning of narrowband THz radiation detectors operating at different frequencies. We also note that within the top candidates, molecules with the same SMILES string were generated multiple times with different 3D structures in the different biasing iterations. As the SCScore and KRR-predicted P values depend only on 2D information, they remain the same for different conformers. However, the PaiNN-predicted P values for raw generated structures and DFT-calculated P values for structures that have undergone geometry optimization can differ, as shown in section S10 and Fig. S24 of the ESI. This further highlights the benefits of working with property predictors that are based on 3D descriptors.
@@
Of the 34 molecules listed in Table SVI, only one compound (generated three times as different conformers, all sharing the same SMILES string) was identified in the PubChem [60, 61] database, Nc1cc(S)c(cc1N)N, which corresponds to 2,4,5-triaminobenzenethiol (Compound Identifier 67981805 [66]). The remaining 31 molecules were not found in PubChem, likely representing novel candidate structures for THz upconversion applications.
@@ IV Conclusions and Outlook
V Conflicts of interest
There are no conflicts to declare.
VI Data availability
Data for this article, including molecular databases in ASE database format, DFT-optimized best candidate molecules, and ASE databases for xTB calculations are available online: https://doi.org/10.6084/m9.figshare.28539995.v3 [67]. The repository also contains the trained ML models, and Jupyter Notebooks and scripts associated with the generation, prediction, and analysis workflows described in this work. Code for the extraction of bonding features from molecular databases and obtaining the principal components of the structural/bonding descriptors has been released in our GSchNetTools package, available at https://github.com/maurergroup/GSchNetTools. The SCScore model used in this work is publicly available at https://github.com/connorcoley/scscore, and files pertaining to retrosynthetic planning with AiZynthFinder are publicly available at https://figshare.com/articles/dataset/AiZynthFinder_a_fast_robust_and_flexible_open-source_software_for_retrosynthetic_planning/12334577.
VII Acknowledgements